1) given mu = 20 and sigma = 3

a) probability(x<18)

b)probability(x>21)

2) given mu = 50 and sigma = 4

a) find 90% confidence interval

b) find 94%
Answer #1

Mแ28 0 Sigma = 3 M-20 ZX-A a) p(x-18) = P(X-4 < 8th - = Plze 18-2 = 8/22 -0.67) I p(x-18) = 02514 6) p(x221) = P(X-4 21-24).b qul - 290.06 2 = 0.03 plZc1.88) ł 0.97 - Z =1.88 %=M+02 = 50 I 4/1.88) = 50 I 7.59 (42.48 , 57.52] १५/ CT 42.48 to 57:59